A population consists of the numbers 5, 20, 9, 4, and 2 with sample size of 3



1
Expert's answer
2022-06-07T04:54:27-0400

We have population values 2,4,5,9,20, population size N=5 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{2+4+5+9+20}{5}=8"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{N}=\dfrac{36+16+9+1+144}{5}=41.2""\sigma=\sqrt{\sigma^2}=\sqrt{41.2}\approx6.4187"


The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_3=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 2,4,5 & 11/3 \\
 \hdashline
 2 & 2,4,9 & 15/3 \\
 \hdashline
 3 & 2,4,20 & 26/3\\
 \hdashline
 4 & 2,5,9 & 16/3 \\
 \hdashline
 5 & 2,5,20 & 27/3 \\
 \hdashline
 6 & 2,9,20 & 31/3 \\
 \hdashline
 7 & 4,5,9 & 18/3 \\
 \hdashline
 8 & 4,5,20 & 29/3 \\
 \hdashline
 9 & 4,9,20 & 33/3 \\
 \hdashline
 10 & 5,9,20 & 34/3 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X})& \bar{X}^2f(\bar{X}) 
\\ \hline
11/3 & 1/10 & 11/30 & 121/90 \\
 \hdashline
 15/3 & 1/10 & 15/30 & 225/90 \\
 \hdashline
 16/3 & 1/10 & 16/30 & 256/90 \\
 \hdashline
 18/3 & 1/10 & 18/30 & 324/90 \\
 \hdashline
 26/3 & 1/10 & 26/30 & 676/90 \\
 \hdashline
 27/3 & 1/10 & 27/30 & 729/90 \\
 \hdashline
 29/3 & 1/10 & 29/30 & 841/90 \\
 \hdashline
 31/3 & 1/10 & 31/30 & 961/90 \\
 \hdashline
 33/3 & 1/10 & 33/30 & 1089/90 \\
 \hdashline
34/3 & 1/10 & 34/30 & 1156/90 \\
 \hdashline
\end{array}"


Mean of sampling distribution 


"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=8=\mu"



The variance of sampling distribution 


"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{6378}{90}-(8)^2=\dfrac{103}{15}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})""\sigma_{\bar{X}}=\sqrt{\dfrac{103}{15}}\approx2.6204"